OpenStudy (anonymous):

how do i find a vertex if its just x^2+2x instead of x^2+2x+3???

OpenStudy (saifoo.khan):

(-1,-1)

OpenStudy (anonymous):

walk me through it please! i know the y intercept already

OpenStudy (saifoo.khan):

y-int is zero.

OpenStudy (saifoo.khan):

You can use, -b/2a It's a shortcut to find the vertex.

OpenStudy (anonymous):

oh ok!

OpenStudy (anonymous):

i thought that was for the axis of symetry though

OpenStudy (saifoo.khan):

Yes, we can substitute the value we get in this equation to get the value of "y" at that point

OpenStudy (anonymous):

ok so to get the axis of symetry u do -b/2a to get the y intercept u plug in zero and mirror across the axis of semetry and to get the vertex u...plug in the axis of semmetry?

OpenStudy (saifoo.khan):

NO! -b/2a to get the "x cordinate" of the vertex. Then substiute that value of x in the equation to get it's y-point.

OpenStudy (anonymous):

the axis of symmetry is simply the vertical line that passes through the x-coordinate of the vertex. To get the y-coordinate of the vertex, plug in the x-coordinate for x and solve f(x)

OpenStudy (anonymous):

OHHHHHH!

OpenStudy (anonymous):

what is the y intercept for x^2+2x

OpenStudy (anonymous):

set x=0 y=0^2+2(0)=0 y-intercept is the point (0,0)

OpenStudy (anonymous):

k then i plug in y which is zero and get 0 again?

OpenStudy (anonymous):

in this case, yes. you know that x=0 for a y-intercept, it's the y-value that you are solving. Every y-intercept will be in the form (0,y)

OpenStudy (anonymous):

so when i do -b/2a i get the axis of symetry

OpenStudy (anonymous):

x=-b/2a for axis of symmetry. Nothing to do with the y-intercept

OpenStudy (anonymous):

and from there yow do i get the vertex? hopefully this will answer everything

OpenStudy (anonymous):

*how

OpenStudy (anonymous):

the x-coordinate is -b/2a. what is the x-coordinate of the vertex?

OpenStudy (anonymous):

ughhh 0,0 and 0,-2?

OpenStudy (anonymous):

-1!!!

OpenStudy (anonymous):

its on the axis line

OpenStudy (anonymous):

-1...right, now what's the y-value?

OpenStudy (anonymous):

0 i think

OpenStudy (anonymous):

f(x)=x^2+2x f(-1)=???

OpenStudy (anonymous):

oh u plug in the x point! -1!!!! so the vertex is -1 -1!!!!! YAAAAAAAAAAAAAAAAAAAAA!!!!!!!!

OpenStudy (anonymous):

U JUST HELPED ME SOOOOOOOO MUCH!

OpenStudy (anonymous):

glad i could help

OpenStudy (anonymous):

ima fan!

OpenStudy (anonymous):

Hey, roadrunner, just saw that you posted the same questions multiple times. Please only post a question once, then wait patiently. Thanks!

Directrix (directrix):

y=2x^2+1 y=2x^2+1 y = 2 x^2 + 0x + 1 a = 2, b = 0, and c = 1 x-coordinate of vertex of parabola and also through which the axis of symmetry passes. x = -b/(2a) x = -(0)/ [(2)(2)] x = -0 / 4 x = 0 x=0 is the equation of the line of symmetry; it is also the y-axis. Take a look at the graph of the parabola and the axis of symmetry at the link: http://www.wolframalpha.com/input/?i=%28y%3D2x%5E2%2B1%29+and+%28x%3D0%29 A person is never too young to begin learning more mathematics nor too old.